What is the amount of heat required to change 2kg of ice at -15°c to water at 70°c ( C=4200jkg-¹k-¹

1
Expert's answer
2022-02-11T08:48:51-0500

Given:

"m=2\:\rm kg"

"c_{ice}=2100\:\rm J/(kg\cdot K)"

"c_{water}=4200\:\rm J/(kg\cdot K)"

"\lambda=330\:000\:\rm J/(kg)"


The total amount of heat

"Q=Q_1+Q_2+Q_3"

"Q=c_{ice}m(0^\circ C-(-15^\circ C))+m\lambda\\
+c_{water}m(70^\circ C-0^\circ C)"

"Q=2100*2*15+2*330\:000
+4200*2*70\\
=1311000\:\rm J"